# Illuminating Black Hole Shadow with Dark Matter Annihilation

### v2

@@ Abstract
The Event Horizon Telescope (EHT) has revolutionized our ability to study black holes by providing unprecedented spatial resolution and unveiling horizon-scale details. With advancements leading to the next-generation EHT, there is potential to probe even deeper into the black hole’s dark region, especially the inner shadow characterized by low-intensity foreground emissions from the jet, thanks to a significant enhancement in dynamic range by two orders of magnitude. We demonstrate how such enhanced observations could transform supermassive black holes into powerful probes for detecting annihilating dark matter, which can form a dense profile in the vicinity of supermassive black holes, by examining the morphology of the black hole image.
I Introduction
@@
The rapid development of the Very Long Baseline Interferometry (VLBI) technique has enabled extraordinarily high angular resolution in radio astronomy. A notable illustration of this progress is the recent imaging of supermassive black holes (SMBHs) achieved by the Event Horizon Telescope (EHT) Akiyama et al. (2019a, 2022), which unveiled detailed astrophysical information in the strong gravity regions. Anticipated upgrades, like the next-generation EHT (ngEHT), promise further enhancements in angular resolution, dynamic range, and baseline coverage Johnson et al. (2023). These advancements are pivotal not only for astrophysical insights but also for exploring fundamental physics, including testing general relativity, examining black hole (BH) properties, and investigating ultralight new bosons Ayzenberg et al. (2023).
@@
Beyond ultralight bosons, another class of potential new particles, often considered as dark matter (DM) candidates, exists around the GeV mass scale. These include weakly interacting massive particles Silk et al. (2010) and sub-GeV DM Knapen et al. (2017). In regions dominated by SMBHs’ gravitational potential, the distribution of particle-like DM can steeply concentrate towards the BH, resulting in densities significantly higher than those near Earth Gondolo and Silk (1999). Thus, SMBHs could serve as effective detectors for DM particles.
@@
A promising approach for observing the DM spike is through indirect detection methods, like observing photons or cosmic rays resulting from potential DM annihilation Gondolo and Silk (1999). These methods leverage the principle that the annihilation rate, directly proportional to the square of the density, increases significantly in regions of high density, thereby enhancing the production of particles. At the EHT’s millimeter radio band, photons are typically produced via synchrotron radiation, where electrons spiral around magnetic field lines Falcke et al. (2000). Understanding the magnetic field structure near the horizon is essential for predicting radio fluxes from DM annihilation. In this study, we employ the best-fit general relativistic magnetohydrodynamic (GRMHD) simulation, which aligns with EHT’s observations of SMBH M87 Akiyama et al. (2019b, 2021a), specifically the magnetically arrested disk (MAD) model Narayan et al. (2003), to calculate the spectrum of DM annihilated electron-positron pairs near the SMBH and their resulting synchrotron radiation. Our results indicate that the density distribution of these pairs is similar in both equatorial and polar regions, contrasting with the MAD model prediction where electrons predominantly inhabit the disk, with lower densities in the jet region. Therefore, we propose utilizing the morphology of BH images, especially future observations of the inner shadow Chael et al. (2021)—a region delineated by the lensed image of the equatorial horizon—to impose constraints on DM annihilation. These constraints appear significantly more stringent than those derived from the total intensity of the image Lacroix et al. (2017); Yuan et al. (2022).
II Electron-Positron Spectrum from Dark Matter Annihilations Near Supermassive Black Holes
@@
The density of DM within a galaxy is typically highest at its center, and the presence of an SMBH can further sharpen this distribution, leading to an increased density towards the center. When considering an initial DM distribution profile where the mass of the central SMBH is sufficiently low to influence the distribution, the slow accretion of ordinary matter onto the SMBH can lead to an adiabatic distortion of DM phase space by the SMBH’s gravitational potential Gondolo and Silk (1999). For instance, in a standard Navarro-Frenk-White (NFW) profile derived from N-body simulations, which neglect the SMBH, the energy density $\rho(r)\propto r^{-1}$ in the central region, where $r$ represents the distance from the galaxy’s center Navarro et al. (1996). The presence of an SMBH, assuming its adiabatic growth, leads to the formation of a DM spike, with its density scaling as $\propto r^{-7/3}$ at the center and transitioning back to the NFW profile in the outer regions. However, this adiabatic formation of a spike can be mitigated by factors such as stellar interactions Vasiliev and Zelnikov (2008). Furthermore, DM annihilation introduces an upper limit on the density, inversely proportional to $\langle\sigma v\rangle$, the thermally averaged cross-section times the relative velocity of DM particles, which effectively regulates the central spike into a flat core Gondolo and Silk (1999).
@@
Depending on their mass and interaction channels with Standard Model particles, DM can annihilate into a spectrum of final-state particles, notably electrons and positrons in this study, which emit synchrotron radiation in magnetic fields Cirelli et al. (2011). These particles, after being produced near an SMBH, move under the influence of both gravitational and magnetic fields. We have developed a comprehensive framework, detailed in Supplemental Material, to calculate the steady-state electron-positron spectrum after propagation in the complex magnetic field environment surrounding an SMBH. This framework represents an advancement over prior spherical models of DM propagation Regis and Ullio (2008); Lacroix et al. (2014, 2017); Yuan et al. (2022), aiming for a more astrophysically precise description.
@@
The background magnetic field utilized for calculating the propagation of electrons and positrons is derived from a high-spin MAD model Narayan et al. (2003), which is supported by observations of the SMBH M87 by the EHT Akiyama et al. (2019b, 2021a). As illustrated in Fig. 1, with the $z$-axis aligned with the BH’s spin, this model reveals an azimuthal and time-averaged electron-positron density $n_{e}$ concentrated primarily in the disk region, with significantly lower densities in the jet region. We utilize an axisymmetric ansatz to fit the magnetic field configurations in both the disk and jet regions, based on GRMHD simulations for the MAD. We assume that the bulk velocities of the electron/positron plasma resulting from DM annihilation align with the magnetic field lines. In the disk region, their radial components are directed towards the BH, with a magnitude scaling as $(2\,r_{g}/r)^{1/2}$, where $r_{g}$ is the gravitational radius of the BH. In the jet region, we account for the rotation of the magnetic field by considering both the inertial centrifugal potential in a co-rotating frame and the gravity potential, which leads to the identification of a zero-velocity area known as the stagnation surface Pu et al. (2015). The initial spectra of DM-annihilated electron-positron pairs are calculated using the PPPC Cirelli et al. (2011) and MadDM Ambrogi et al. (2019) packages.
@@
Figure 1: Electron-positron number density $n_{e}$ from an azimuthally-averaged MAD model based on GRMHD (left) versus from DM annihilation (right). The $z$-axis is aligned with the BH spin, adopting a dimensionless value of $a_{J}=0.9375$ in the GRMHD simulation. The DM profile assumes a standard spike originating from an initial NFW profile, featuring a flat core at the center due to annihilation. In this example, the DM mass is $m_{\rm DM}=95\,$GeV, with the annihilation channel being $b\bar{b}$ and $\langle\sigma v\rangle=10^{-27}\,\text{cm}^{3}\text{s}^{-1}$.
@@
The right panel of Fig. 1 presents an example of the resulting electron and positron density, excluding electrons from the GRMHD MAD model, where the DM mass is $m_{\rm DM}=95\,$GeV, with the annihilation channel being $b\bar{b}$ and $\langle\sigma v\rangle=10^{-27}\,\text{cm}^{3}\text{s}^{-1}$. The dark blue region, indicative of areas outside the stagnation surface, is not considered for conservative estimations. The densities in the jet and disk regions prove to be comparable, differing markedly from the GRMHD MAD case. This discrepancy arises because the distribution of electrons/positrons is expelled by the jet in the GRMHD scenario, whereas DM annihilation continuously supplies them within the jet cone.
III Constraints from the EHT and ngEHT
@@
Figure 2: Top: normalized intensity map ($I/I_{\rm max}$) in logarithmic scale comparing the MAD model from GRMHD (left) to DM annihilation (right), for the SMBH M87 with $a_{J}=0.9375$ at an inclination angle of $\theta_{o}=163^{\circ}$. Normalization is against the maximum intensity in the MAD model, $I_{\rm max}$. The critical curve and the inner shadow’s boundary are denoted by dashed and green lines, respectively. Both maps exhibit a total intensity of 0.60.6 0.60.6 Jy. Bottom: one-dimensional normalized intensity profile along the axis ($\alpha$), centered on the black hole (BH) and perpendicular to the spin projection, showcasing contrasts between GRMHD MAD (blue) and DM annihilation (red) scenarios. Dark and light shades indicate the original image and its Gaussian-smoothed counterpart with an $1\,r_{g}$ kernel, respectively.
@@
Constraints on DM annihilation require that radiation from this process remains below the astrophysical background, positioning astrophysical radiation as a pivotal factor in determining the exclusion parameter space. We focus on the synchrotron radiation from the MAD model at 230230 230230 GHz radio frequency band. Our analysis concentrates on the synchrotron radiation emanating from the MAD model within the 230230 230230 GHz radio frequency band. By employing the RAPTOR covariant radiative transfer package Bronzwaer et al. (2018), we generate a horizon-scale intensity ($I$) map for MAD outside an SMBH with $a_{J}=0.9375$ and an inclination angle of $\theta_{o}=163^{\circ}$, consistent with observations of M87, depicted in the top left panel of Fig. 2. Subsequent application of Gaussian smearing with an approximately $2\,r_{g}$ kernel to simulate the current EHT’s angular resolution yields an image that closely matches the observed data Akiyama et al. (2019b), with a total flux of approximately 0.60.6 0.60.6 Jy
@@
The BH image, presented on a logarithmic scale, reveals fascinating morphological features that reflect the interplay between Kerr spacetime and the sources of emission. The dashed line in Fig. 2 delineates the critical curve, beyond which null geodesics are captured into bound orbits around the BH, creating a photon ring from light circling multiple times Podurets (1965); Ames and Thorne (1968); Luminet (1979). Consequently, the intensity observed in the adjacent region results from the cumulative effect of photons undergoing multiple loops, achieving their peak in the map Johannsen and Psaltis (2010). Within the critical curve, backward geodesics terminate at the BH horizon, known as the BH ‘shadow’ Falcke et al. (2000). Nonetheless, this area is not completely devoid of light, as emission, proportional to electron density $n_{e}$ and magnetic field strength $B$, reaches up to the horizon. As shown in Fig. 1 and Supplemental Material, the distribution of $B$ slightly changes with the polar angle, whereas $n_{e}$ is predominantly found in the accretion disk, with much lower concentrations in the jet region. This distribution creates a geometrically thin emission profile mainly in the equatorial disk, leading to sharp intensity depression in the central area, identified as the ‘inner shadow’ Chael et al. (2021). The boundary of this inner shadow, marked in green in Fig. 2, outlines the lensed contour of the BH’s equatorial horizon.
@@
The current EHT’s capability to detect the inner shadow is hindered by its limited dynamic range ($\sim 10$) and angular resolution on the intensity dip. In contrast, the ngEHT is anticipated to achieve a substantial enhancement in both dynamic range ($\sim 1000$) and angular resolution, equivalent to a Gaussian kernel of approximately $1\,r_{g}$. This improvement holds the promise of capturing this fascinating feature and delving into the BH horizon Chael et al. (2021). The bottom panel of Fig. 2 displays the one-dimensional intensity distribution along an axis that crosses the BH center and is perpendicular to the projection of its spin. The use of dark and light shades serves to distinguish between the original image and the image after applying Gaussian smearing. Notably, within the inner shadow, the jet’s foreground emission is markedly lower—by two orders of magnitude—than the peak intensity, falling well within the ngEHT’s enhanced dynamic range.
@@
The pronounced dimming within the inner shadow establishes a solid framework for testing DM annihilation. As shown in Fig. 1, the electron-positron density $n_{e}$, induced by DM annihilation, attains comparable levels in both the jet and disk regions, akin to the results observed in spherical emission models Vincent et al. (2022). This morphology allows the DM-induced density to exceed the jet’s emission in the inner shadow area. The intensity map, derived solely from DM annihilation and presented in the top right panel of Fig. 2 with the benchmark parameters from Fig. 1, reveals a total intensity nearly matching the astrophysical emission of 0.60.6 0.6\,0.6Jy. In contrast, the inner shadow is significantly illuminated, lacking the expected intensity reduction. This discrepancy becomes even more evident in the bottom panel of Fig. 2, where the red line for DM annihilation markedly exceeds the expected intensity levels within the inner shadow, deviating from the astrophysical emission patterns.
@@
Figure 3: Constraints on DM annihilation into $b\bar{b}$ (top panel) and $e^{+}e^{-}$ (bottom panel) channels, incorporating total intensity constraints observed by the EHT in blue, and constraints inferred from BH image morphology by the EHT (green) and the ngEHT (black). The shaded region within the ngEHT exclusion indicates the ratio of local intensity between DM annihilation and astrophysical emissions from the MAD model in the inner shadow. The thermal relic annihilation cross-section Steigman et al. (2012) is depicted by the purple line. Previous constraints from various sources, including Planck CMB observations Slatyer (2016), Fermi-LAT Ackermann et al. (2015), H.E.S.S. Abdallah et al. (2016), AMS Bergstrom et al. (2013), and XMM-Newton Cirelli et al. (2023); De la Torre Luque et al. (2023), as well as spike-based constraints Lacroix et al. (2015); Balaji et al. (2023), are shown for comparative analysis in dashed and dot-dashed lines, respectively.
@@
In Fig. 3, we illustrate the exclusion regions derived from analyses using both the EHT and the ngEHT, targeting DM annihilation channels into $b\bar{b}$ and $e^{+}e^{-}$, associated with a DM spike around M87. The constraint, illustrated in blue and based on the total intensity observed by the EHT, is derived by ensuring that the total intensity from DM annihilation does not exceed the overall astrophysical emission. Both are considered over a region extending up to $10\,r_{g}$ from the SMBH. Conversely, the prospective morphological constraint from the ngEHT, depicted in black, imposes a stricter criterion by demanding that the local intensity resulting from DM annihilation consistently remains below the astrophysical emission within the region extending to $10\,r_{g}$. The exclusion area defined by the ngEHT’s criteria markedly expands beyond that established by the EHT, underscoring the ngEHT’s increased sensitivity to the BH image’s morphology. The density within the exclusion zone represents the ratio of local intensity between DM annihilation and MAD-induced emissions in the inner shadow region. Notably, the ngEHT’s exclusion region primarily results from the enhanced visibility of the inner shadow due to DM annihilation. However, in cases with a high annihilation cross-section, the surpassing of astrophysical emission by DM annihilation becomes evident in the outer regions of the observer plane. This is due to the formation of a core of uniform decnsity by the annihilating dark matter. The EHT’s morphological constraint, highlighted in green, applies criteria similar to those used by the ngEHT but includes an additional condition. Specifically, after Gaussian smearing with a $2\,r_{g}$ kernel, the local intensity resulting from DM annihilation must exceed the dynamic range threshold of $1/10$. In the $b\bar{b}$ channel, the exclusion region corresponds closely to areas exhibiting the highest electron-positron densities as shown in Supplemental Material. Conversely, the exclusion criteria for the $e^{+}e^{-}$ channel exhibit a notable turnover at around 0.10.1 0.10.1 GeV. This shift is attributed to the inadequacy of synchrotron radiation at 230230 230230 GHz for $m_{\rm DM}$ below 0.10.1 0.10.1 GeV.
@@
The purple line in Fig. 3 indicates the theoretically well-motivated thermal relic annihilation cross-section Steigman et al. (2012), which the ngEHT could potentially probe for DM masses up to approximately a TeV. For context, the figure also presents previous constraints obtained from the Planck observation of the cosmic microwave background (CMB) Slatyer (2016), as well as constraints from the Fermi Large Area Telescope (Fermi-LAT) Ackermann et al. (2015), the High Energy Stereoscopic System (H.E.S.S.) Abdallah et al. (2016), the Alpha Magnetic Spectrometer Experiment (AMS) Bergstrom et al. (2013), and the X-ray Multi-Mirror Mission (XMM-Newton) Cirelli et al. (2023); De la Torre Luque et al. (2023), depicted with dashed lines. Additionally, dot-dashed lines highlight previous constraints that are specific to a DM spike profile Lacroix et al. (2015); Balaji et al. (2023). Notably, some previous constraints derived from synchrotron radiation may lead to an overestimation of the exclusion region, attributed to the adoption of a simplified magnetic field profile Lacroix et al. (2015).
IV Discussion

### v3

@@ Abstract
The Event Horizon Telescope (EHT) has significantly advanced our ability to study black holes, achieving unprecedented spatial resolution and revealing horizon-scale structures. Notably, these observations feature a distinctive dark shadow–primarily arising from faint jet emissions–surrounded by a bright photon ring. Anticipated upgrades of the EHT promise substantial improvements in dynamic range, enabling deeper exploration of low-background regions, particularly the inner shadow defined by the lensed equatorial horizon. Our analysis shows that observations of these regions transform supermassive black holes into powerful probes for annihilating dark matter, which is expected to accumulate densely in their vicinity. By analyzing the black hole image morphology and performing electron-positron propagation calculations in realistic plasma backgrounds derived from general relativistic magnetohydrodynamic simulations, we set stringent constraints on dark matter annihilation, requiring contributions below the astrophysical emission. These constraints, derived from both current EHT observations and projections for future upgraded arrays, exclude a substantial region of previously unexplored parameter space and remain robust against astrophysical uncertainties, including black hole spin and plasma temperature variations.
@@
Introduction–The rapid development of the very long baseline interferometry (VLBI) technique has enabled extraordinarily high angular resolution in radio astronomy. A notable illustration of this progress is the recent imaging of supermassive black holes (SMBHs) achieved by the Event Horizon Telescope (EHT) Akiyama et al. (2019a, 2022), which unveiled detailed astrophysical information in the strong gravity regions. Anticipated future upgrades of the EHT, including possible ground-based expansions and the proposed space-based Black Hole Explorer (BHEX), promise further enhancements in angular resolution, dynamic range, and baseline coverage Johnson et al. (2023, 2024). These advancements are pivotal not only for astrophysical insights but also for exploring fundamental physics, including testing general relativity, examining black hole (BH) properties, and investigating new fundamental fields Ayzenberg et al. (2025).
@@
One class of potential new particles, often considered as dark matter (DM) candidates, is found around the GeV mass scale. These candidates include weakly interacting massive particles (WIMPs) Silk et al. (2010) and sub-GeV DM Knapen et al. (2017). In regions dominated by SMBHs’ gravitational potential, the distribution of particlelike DM can steeply concentrate toward the BH, resulting in densities significantly higher than those near Earth Gondolo and Silk (1999). Thus, SMBHs could serve as effective detectors for DM particles.
@@
A promising approach for observing the DM spike is through indirect detection methods, like observing photons or cosmic rays resulting from potential DM annihilation Gondolo and Silk (1999). These methods leverage the principle that the annihilation rate, directly proportional to the square of the density, increases significantly in regions of high density, thereby enhancing the production of particles. At the EHT’s millimeter radio band, photons are typically produced via synchrotron radiation, where electrons spiral around magnetic field lines Falcke et al. (2000). Understanding the magnetic field structure near the horizon is essential for predicting radio fluxes from DM annihilation. In this study, we employ the best-fit general relativistic magnetohydrodynamic (GRMHD) simulation, which aligns with EHT’s observations of SMBH M87 Akiyama et al. (2019b, 2021a), specifically the magnetically arrested disk (MAD) model Narayan et al. (2003), to calculate the spectrum of DM annihilated electron-positron pairs near the SMBH and their resulting synchrotron radiation. Our results indicate that the density distribution of these pairs is similar in both equatorial and polar regions, contrasting with the MAD model prediction where electrons predominantly inhabit the disk, with lower densities in the jet region. Therefore, we utilize the morphology of BH images to impose constraints on DM annihilation, analyzing both current observations of the dark region resulting from faint jet emissions and future observations of the inner shadow–a region delineated by the lensed image of the equatorial horizon Chael et al. (2021). These constraints, based on the morphology, prove significantly more stringent than those derived from the total intensity of the image Lacroix et al. (2017); Yuan et al. (2022).
@@
Electron-Positron Spectrum from Dark Matter Annihilations Near Supermassive Black Holes–The density of DM within a galaxy is typically highest at its center, and the presence of an SMBH can further sharpen this distribution, leading to an increased density toward the center. When considering an initial DM distribution profile where the mass of the central SMBH is sufficiently low to influence the distribution, the slow accretion of ordinary matter onto the SMBH can lead to an adiabatic distortion of DM phase space by the SMBH’s gravitational potential Gondolo and Silk (1999). For instance, in a standard Navarro-Frenk-White (NFW) profile derived from N-body simulations, which neglect the SMBH, the energy density $\rho(r)\propto r^{-1}$ in the central region, where $r$ represents the distance from the galaxy’s center Navarro et al. (1996). The presence of an SMBH, assuming its adiabatic growth, leads to the formation of a DM spike, with its density scaling as $\propto r^{-7/3}$ at the center and transitioning back to the NFW profile in the outer regions. However, this adiabatic formation of a spike can be mitigated by factors such as stellar interactions Vasiliev and Zelnikov (2008). Furthermore, DM annihilation introduces an upper limit on the density, inversely proportional to $\langle\sigma v\rangle$, the thermally averaged cross section times the relative velocity of DM particles, which effectively regulates the central spike into a flat core Gondolo and Silk (1999).
@@
Depending on their mass and interaction channels with standard model particles, DM can annihilate into a spectrum of final-state particles, notably electrons and positrons in this study, which emit synchrotron radiation in magnetic fields Cirelli et al. (2011). These particles, after being produced near an SMBH, move under the influence of both gravitational and magnetic fields. We have developed a comprehensive framework, detailed in Supplemental Material Sup , to calculate the steady-state electron-positron spectrum after propagation in the complex magnetic field environment surrounding an SMBH. We also include a discussion of the approximations employed, such as the omission of certain general relativistic effects, which have only a minor impact on our results. This framework represents an advancement over prior spherical models of DM propagation Regis and Ullio (2008); Lacroix et al. (2014, 2017); Yuan et al. (2022), aiming for a more astrophysically precise description.
@@
The background magnetic field utilized for calculating the propagation of electrons and positrons is derived from MAD-type models Narayan et al. (2003), which is supported by observations of the SMBH M87 by the EHT Akiyama et al. (2019b, 2021a). In Fig. 1, we show an example of the azimuthally and time-averaged electron and positron density distribution, $n_{e}$, for a BH spin of $a_{J}=0.9375$. The $z$ axis is aligned with the BH’s spin. The density is primarily concentrated in the disk region, with significantly lower values in the jet region. We utilize an axisymmetric ansatz to fit the magnetic field configurations in both the disk and jet regions, based on GRMHD simulations for the MAD. We assume that the bulk velocities of the electron and positron plasma resulting from DM annihilation align with the magnetic field lines. In the disk region, their radial components are directed toward the BH, with a magnitude scaling as $(2\,r_{g}/r)^{1/2}$, where $r_{g}$ is the gravitational radius of the BH. In the jet region, we account for the rotation of the magnetic field by incorporating both the inertial centrifugal potential in the corotating frame and the gravitational potential, leading to the identification of a zero-velocity surface known as the stagnation surface Pu et al. (2015, 2017); Pu and Takahashi (2020). The initial spectra of DM annihilated electron and positron pairs are calculated using the PPPC Cirelli et al. (2011) and MadDM Ambrogi et al. (2019) packages.
@@
Figure 1: Electron-positron number density $n_{e}$ from an azimuthally averaged MAD model based on GRMHD (left) versus from DM annihilation (right). The $z$ axis is aligned with the BH spin, adopting a BH spin of $a_{J}=0.9375$ in the GRMHD simulation. The DM profile assumes a standard spike originating from an initial NFW profile, with a DM mass of $m_{\rm DM}\!=\!240\,\mathrm{GeV}$, annihilation via the $b\bar{b}$ channel, and a cross section of $\langle\sigma v\rangle=10^{-27}\,\mathrm{cm}^{3}\,\mathrm{s}^{-1}$. The black dashed line marks the jet boundary, while the gray dashed line indicates the stagnation surface.
@@
The right panel of Fig. 1 presents an example of the resulting electron and positron density, excluding electrons from the GRMHD MAD model, where the DM mass is $m_{\rm DM}\!=\!240\,$GeV, with the annihilation channel being $b\bar{b}$ and $\langle\sigma v\rangle=10^{-27}\,\text{cm}^{3}\text{s}^{-1}$. The black dashed line marks the jet boundary, while the gray dashed line indicates the stagnation surface. The densities in the jet and disk regions prove to be comparable, differing markedly from the GRMHD MAD case. This discrepancy arises because the distribution of electrons and positrons is expelled by the jet in the GRMHD scenario, whereas DM annihilation continuously supplies them within the jet cone.
@@
Figure 2: Top: normalized intensity map ($I/I_{\rm max}$) in logarithmic scale comparing the MAD model from GRMHD (left) to DM annihilation (right), for the SMBH M87 with $a_{J}=0.9375$ at an inclination angle of $\theta_{o}=163^{\circ}$. Normalization is against the maximum intensity in the MAD model, $I_{\rm max}$. The critical curve and the inner shadow’s boundary are denoted by dashed and green lines, respectively. Both maps exhibit a total intensity of $0.6$ Jy. Bottom: one-dimensional normalized intensity profile along the axis ($\alpha$), centered on the BH and perpendicular to the spin projection, showcasing contrasts between GRMHD MAD (blue) and DM annihilation (red) scenarios. Dark and light shades indicate the original image and its Gaussian-smoothed counterpart with an $1\,r_{g}$ kernel, respectively.
@@
Constraints from the EHT and future EHT upgrades–Constraints on DM annihilation require that radiation from this process remains below the astrophysical background, positioning astrophysical radiation as a pivotal factor in determining the exclusion parameter space. We focus on the synchrotron radiation from the MAD model at $230$ GHz radio frequency band. Our analysis concentrates on the synchrotron radiation emanating from the MAD model within the $230$ GHz radio frequency band. By employing the RAPTOR covariant radiative transfer package Bronzwaer et al. (2018), we generate a horizon-scale intensity ($I$) map for MAD outside an SMBH with an inclination angle of $\theta_{o}=163^{\circ}$, consistent with observations of M87, depicted in the top left panel of Fig. 2. Subsequent application of Gaussian smearing with an approximately $2\,r_{g}$ kernel to simulate the current EHT’s angular resolution yields an image that closely matches the observed data Akiyama et al. (2019b), with a total flux of approximately $0.6$ Jy.
@@
The BH image, presented on a logarithmic scale, reveals fascinating morphological features that reflect the interplay between Kerr spacetime and the sources of emission. The dashed line in Fig. 2 delineates the critical curve, beyond which null geodesics are captured into bound orbits around the BH, creating a photon ring from light circling multiple times Podurets (1965); Ames and Thorne (1968); Luminet (1979). Consequently, the intensity observed in the adjacent region results from the cumulative effect of photons undergoing multiple loops, achieving their peak in the map Johannsen and Psaltis (2010). Within the critical curve, backward geodesics terminate at the BH horizon, known as the BH “shadow” Falcke et al. (2000). Nonetheless, this area is not completely devoid of light, as emission, proportional to electron density $n_{e}$ and magnetic field strength $B$, reaches up to the horizon. As shown in Fig. 1 and Supplemental Material, the distribution of $B$ slightly changes with the polar angle, whereas $n_{e}$ is predominantly found in the accretion disk, with much lower concentrations in the jet region. This distribution creates a geometrically thin emission profile mainly in the equatorial disk, leading to sharp intensity depression in the central area, identified as the “inner shadow” Chael et al. (2021). The boundary of this inner shadow, marked in green in Fig. 2, outlines the lensed contour of the BH’s equatorial horizon.
@@
The current EHT’s capability to detect the inner shadow is hindered by its limited dynamic range ($\sim 10$) and angular resolution on the intensity dip. In contrast, future EHT upgrades are anticipated to achieve a substantial enhancement in both dynamic range ($\sim 1000$) and angular resolution, equivalent to a Gaussian kernel of approximately $1\,r_{g}$. This improvement holds the promise of capturing this fascinating feature and delving into the BH horizon Chael et al. (2021). The bottom panel of Fig. 2 displays the one-dimensional intensity distribution along an axis that crosses the BH center and is perpendicular to the projection of its spin. The use of dark and light shades serves to distinguish between the original image and the image after applying Gaussian smearing. Notably, within the inner shadow, the foreground emission from the jet region is significantly suppressed–by roughly 2 orders of magnitude compared to the peak intensity–and, thus, well within the detection capabilities of the proposed EHT upgrades Chael et al. (2021).
@@
The pronounced dimming within the shadow region establishes a solid framework for testing DM annihilation. As shown in Fig. 1, the electron-positron density $n_{e}$, induced by DM annihilation, attains comparable levels in both the jet and disk regions, akin to the results observed in spherical emission models Vincent et al. (2022). This morphology allows the DM-induced density to exceed the jet’s emission in the shadow area. The intensity map, derived solely from DM annihilation and presented in the top right panel of Fig. 2 with the benchmark parameters from Fig. 1, reveals a total intensity nearly matching the astrophysical emission of $0.6\,$Jy. In contrast, the inner shadow is significantly illuminated, lacking the expected intensity reduction. This discrepancy becomes even more evident in the bottom panel of Fig. 2, where the red line for DM annihilation markedly exceeds the expected intensity levels within the inner shadow, deviating from the astrophysical emission patterns.
@@
Figure 3: Constraints on DM annihilation channels $b\bar{b}$ (top) and $e^{+}e^{-}$ (bottom), derived from current EHT observations (diagonal stripes), which include criteria based on total intensity (blue) and BH image morphology (green), as well as projections for the EHT upgrades (black). The pink shades illustrate the ratio of local intensity between DM annihilation and astrophysical emissions from the MAD model in the inner shadow. The thermal relic annihilation cross section Steigman et al. (2012) is depicted by the purple line. Previous constraints from various sources, including Planck CMB observations Slatyer (2016), Fermi-LAT Ackermann et al. (2015), H.E.S.S. Abdallah et al. (2016), AMS Bergstrom et al. (2013), and XMM-Newton Cirelli et al. (2023); De la Torre Luque et al. (2023), as well as spike-based constraints Lacroix et al. (2015); Balaji et al. (2023), are shown for comparative analysis in dashed and dot-dashed lines, respectively.
@@
In Fig. 3, we illustrate the exclusion regions derived from analyses using both the current EHT observations and the future prospects, targeting DM annihilation channels $b\bar{b}$ and $e^{+}e^{-}$, associated with a DM spike around M87. The regions with diagonal stripes represent constraints from current EHT observations. The regions enclosed by the blue lines represent exclusion constraints, ensuring that the total intensity from DM annihilation does not exceed the overall astrophysical emission within the EHT field of view, which extends to approximately $16\,r_{g}$ from the BH. The morphological constraints, shown within the green lines, impose a stricter criterion by demanding that the local intensity from DM annihilation, after Gaussian smearing with a $2\,r_{g}$ kernel, consistently remains below the astrophysical emission within $10\,r_{g}$ and exceeds the dynamic range threshold of $1/10$. The prospective morphological constraints from the EHT upgrades, depicted in black, significantly expand beyond that established by the current EHT, underscoring the increased sensitivity to the BH image’s morphology. The density within the exclusion zone represents the ratio of local intensity between DM annihilation and MAD-induced emissions in the inner shadow region. The EHT upgrades’ exclusion region primarily arises from the enhanced dynamic range of approximately $1000$. In the $b\bar{b}$ channel, the exclusion region closely aligns with areas exhibiting the highest electron-positron densities as shown in Supplemental Material. In contrast, the exclusion contour for the $e^{+}e^{-}$ channel exhibits a notable turnover near $0.1$ GeV, due to the reduced efficiency of synchrotron radiation at $230$ GHz for $m_{\rm DM}$ below this threshold.
Several astrophysical uncertainties can influence our results, even within the MAD model class favored by multiple EHT observations of M87 Akiyama et al. (2019c, 2021b, 2023). One such parameter is the BH spin $a_{J}$, which is currently favored to be high Akiyama et al. (2019c), but has only a lower bound of $a_{J}\geq 0.5$ Cruz-Osorio et al. (2022). While the main text presents results using $a_{J}=0.9375$, we also perform a full analysis for $a_{J}=0.5$. The resulting constraints, shown in Supplemental Material, vary only slightly from the fiducial case, demonstrating the robustness of our conclusions. We additionally explore variations in plasma-related parameters such as $R_{\rm high}$ and $R_{\rm low}$, which characterize the proton-to-electron temperature ratio in regions of weak and strong magnetization, respectively Moscibrodzka et al. (2016); Bronzwaer et al. (2018). These variations likewise show minimal impact on our exclusion limits.
@@
The purple line in Fig. 3 denotes the theoretically well-motivated thermal relic annihilation cross section Steigman et al. (2012), which the EHT upgrades could potentially probe for DM masses up to approximately $10$ TeV. For context, the figure also presents previous constraints obtained from the Planck observation of the cosmic microwave background (CMB) Slatyer (2016), as well as constraints from the Fermi Large Area Telescope (Fermi-LAT) Ackermann et al. (2015), the High Energy Stereoscopic System (H.E.S.S.) Abdallah et al. (2016), the Alpha Magnetic Spectrometer Experiment (AMS) Bergstrom et al. (2013), and the X-ray Multi-Mirror Mission (XMM-Newton) Cirelli et al. (2023); De la Torre Luque et al. (2023), depicted with dashed lines. Additionally, dot-dashed lines highlight previous constraints that are specific to a DM spike profile Lacroix et al. (2015); Balaji et al. (2023). Notably, some previous constraints derived from synchrotron radiation may lead to an overestimation of the exclusion region, attributed to the adoption of a simplified magnetic field profile Lacroix et al. (2015). Our constraints improve upon these existing bounds primarily due to the EHT’s superior angular resolution and dynamic range, which enable probing DM annihilation in a compact region with minimal astrophysical foreground contamination.